OpenStudy (anonymous):

Elena is thinking about putting $200 in a savings account that earns 4% interest compounded semiannually. She wants to keep that money in the account for 4 years. Which of the formulas below can help her calculate how much money she will have at the end of the 4 years? A. $200(1 + 0.04)4 B. $200(1 + 0.02)8 C. $200(1 + 0.02)4 D. $200(1 + 0.08)2

OpenStudy (anonymous):

A=P(1+R/100)^t here P=200$, R=4% annual=2% semiannual t=4 years=8 semiannual A=200(1+2/100)^8=200(1+.02)^8